Consider a wheel with a 45 cm radius that is constantly spinning at 30 rpm. Determine the linear and angular speeds for points at distances 15 cm, 30 cm and 45 cm from the center of rotation.
Answer #1

here,

the angular speed , w = 30 rpm

the angular speed is constant , w = 30 rpm = 3.14 rad/s

for r1 = 15 cm = 0.15 m

the linear speed , v1 = w * r1

v1 = 3.14 * 0.15 m/s = 0.471 m/s

for r2 = 30 cm = 0.3 m

the linear speed , v2 = w * r2

v2 = 3.14 * 0.3 m/s = 0.942 m/s

for r3 = 45 cm = 0.45 m

the linear speed , v3 = w * r3

v1 = 3.14 * 0.45 m/s = 1.413 m/s